Suppose ypou make $65,000 a year. You want to purchase a $300,000 house with a 90% LTV loan. The current 30-year FRM interest rate is 6.5%. Your monthly insurance and property tax payment add up to $250. The lender allows a maximum Total Housing Expenses to Income rate of 35%. Will you qualify for this loan? (using financial calculator)
Answer #1

Using financial calculator
N=12*30
I/Y=6.5%/12
FV=0
PV=-300000*90%=-270000
CPT PMT=$1,706.5837

Total monthly house expenses=1706.5837+250=$1,956.5837

Monthly housing expenses as % of monthly income=1956.5837/(65000/12)=36.122%

As this is more than 35%, one would not qualify for the loan